Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates:

Machine-hours required to support estimated production 154,000
Fixed manufacturing overhead cost $ 652,000
Variable manufacturing overhead cost per machine-hour $ 4.20

Required:

1. Compute the plantwide predetermined overhead rate.

2. During the year, Job 400 was started and completed. The following information was available with respect to this job:

Direct materials $ 330
Direct labor cost $ 300
Machine-hours used 33

Compute the total manufacturing cost assigned to Job 400.

3. If Job 400 includes 60 units, what is the unit product cost for this job?

4. If Moody uses a markup percentage of 120% of its total manufacturing cost, then what selling price per unit would it have established for Job 400?
Answer #1

1.

Plantwide predetermined overhead = ($652,000 / 154,000) + $4.20 = $8.43

2.

Total manufacturing cost = $330 + $300 + (33 X $8.43) = $908.19

3.

Unit product cost = $908.19 / 60 = $15.14

4.

Selling price per unit = $15.14 + 120% = $33.31